Please use the information below for the following questions:

A group of 25 people signed up for a weight-loss class. Each individual was asked why they wanted to lose weight. The choices were: H (health reasons); C (cosmetic reasons); or O (other). The results were as follows:

H H C H O C C H C O O H C H H C H H O H H O C H C


a. Determine the frequency for health ( ), cosmetic( ) , and other ( )

b. Determine the relative frequency (compute to two decimal places) for health( ), cosmetic( ) , and other( )  

c. Determine the frequency percentage for health( ) %, cosmetic( ) %, and other( ) %
Answer #1

a) Number of health = 12

Number of cosmetic= 8

Number of other =5

b)Relative frequency of helath = 12/25 = 0.48

Cosmetic =8/25= 0.32

Other= 5/25 = 0.2

c) Frequency percentage of health = 12/25 *100 =48%

Cosmetic = 8/25*100= 32%

Other =5/25*100 =20%